OpenStudy (anonymous):

\[|Z+ \frac{ 9 }{ Z }| = 6\] ,Then find the Greatest value of |Z| ?

OpenStudy (anonymous):

\[\left| \frac{Z^2+9}{Z} \right|=6\] top part always positive \[Z^2+9=6\left| Z \right|\] \[Z^2-6\left| Z \right|+9=(Z-3)(Z-3)\]

OpenStudy (anonymous):

\[(Z-3)^2=0\] Z=3 since Z has no other valueZ=3

OpenStudy (anonymous):

Lol....i tried this way.......the answer shuld be 3 + sqrt (18)

OpenStudy (anonymous):

since\[\left| Z \right|=\sqrt{Z^2}=a\] \[Z^2+6\sqrt{Z^2}+9=0\] \[a-6a^2+9=0\] \[6a^2-a-9=0\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i made a wrong substitution\[a^2-6a+9=0\] \[a=3\] \[Z=\pm3\]

OpenStudy (anonymous):

using calculus method we derive\[Z^2+9=6\left| Z \right|\] \[2Z=6\frac{Z}{|Z|}\] \[|Z|=3\]

OpenStudy (shubhamsrg):

Isn't answer 27 ?

OpenStudy (anonymous):

the answer shuld be 3 + sqrt (18)

OpenStudy (shubhamsrg):

ohh I mean sqrt27 hmm, still now getting.

OpenStudy (shubhamsrg):

not*

OpenStudy (anonymous):

ok...any how Can u show Ur Steps ?

OpenStudy (shubhamsrg):

I got messed up. Sorry Gimme some time.

OpenStudy (anonymous):

@mukushla

OpenStudy (shubhamsrg):

Let z=x+iy | (x+iy)^2 + 9| = 6sqty(x^2 + y^2) |(x^2 -y^2 + 9) + i(2xy)| = 6sqrt(x^2 + y^2) =>(x^2 -y^2 + 9)^2 + (2xy)^2 = 36(x^2 + y^2) We need to maximize (x^2 + y^2) with the given info that (x^2 -y^2 + 9)^2 + (2xy)^2 = 36(x^2 + y^2) Yes the answer is 3 + sqrt18